# PTR: Precision-Driven Tool Recommendation for Large Language Models

### v1

@@ Abstract
By augmenting Large Language Models (LLMs) with external tools, their capacity to solve complex problems has been significantly enhanced. However, despite ongoing advancements in the parsing capabilities of LLMs, incorporating all available tools simultaneously in the prompt remains impractical due to the vast number of external tools. Consequently, it is essential to provide LLMs with a precise set of tools tailored to the specific task, considering both quantity and quality. Current tool retrieval methods primarily focus on refining the ranking list of tools and directly packaging a fixed number of top-ranked tools as the tool set. However, these approaches often fail to equip LLMs with the optimal set of tools prior to execution, since the optimal number of tools for different tasks could be different, resulting in inefficiencies such as redundant or unsuitable tools, which impede immediate access to the most relevant tools. This paper addresses the challenge of recommending precise toolsets for LLMs. We introduce the problem of tool recommendation, define its scope, and propose a novel Precision-driven Tool Recommendation (PTR) approach. PTR captures an initial, concise set of tools by leveraging historical tool bundle usage and dynamically adjusts the tool set by performing tool matching, culminating in a multi-view-based tool addition. Additionally, we present a new dataset, RecTools, and a metric, TRACC, designed to evaluate the effectiveness of tool recommendation for LLMs. We further validate our design choices through comprehensive experiments, demonstrating promising accuracy across two open benchmarks and our RecTools dataset.
@@ 1 Introduction
Large Language Models (LLMs) have established themselves as powerful intermediaries, demonstrating remarkable impacts across a variety of downstream tasks, including text generation, code debugging, and personalized recommendations (Brown et al., 2020; Touvron et al., 2023; Nam et al., 2024; Chen et al., 2024; Zhao et al., 2024). However, as these models continue to evolve, they still struggle to solve highly complex problems due to limitations arising from their pre-training data (Mialon et al., 2023; Mallen et al., 2022; Yuan et al., 2023). To expand the potential of LLMs in managing more complex tasks efficiently, recommendations at various levels have been increasingly applied to LLMs. Typically, memory recommendations (Borgeaud et al., 2022) and knowledge-based recommendations (Gao et al., 2023; Hu et al., 2023) enhance consistency and context awareness in ongoing tasks for LLMs, while data augmentation recommendations (Xu et al., 2020) facilitate the inclusion of additional data to augment training. Furthermore, architecture recommendations (Elsken et al., 2019; Fedus et al., 2022) and prompt recommendations (Shin et al., 2020; Pryzant et al., 2023; Liu et al., 2023) optimize efficiency and generate more relevant outputs. Simultaneously, to reduce the cognitive load on LLMs and enhance their complex problem-solving capabilities by enabling actions beyond natural language processing, it is crucial to augment LLMs with recommendations of optimal external tool sets, an aspect currently lacking in existing recommendation frameworks for LLMs. Furthermore, this approach will be helpful to address the challenge of input length limitations encountered when incorporating a large number of external tools into the prompt. Providing LLMs with a precise and dynamically adaptable recommended toolset can help to enhance the effectiveness of LLM’s task-solving ability.
@@
Considering that the capability of LLMs to master and control external tools is instrumental in overcoming some of their fundamental weaknesses, the field of tool retrieval—which aims to identify the top-K most suitable tools for a given query from a vast set of tools—has been increasingly explored. The advent of tool retrieval (Zhuang et al., 2023; Li et al., 2023; Tang et al., 2023; Yang et al., 2024) signifies a nuanced evolution, most directly employing term-based methods (Sparck Jones, 1972; Robertson et al., 2009) or semantic-based techniques (Kong et al., 2023; Yuan et al., 2024; Gao et al., 2024). Generally, the primary objective of these methods is to refine the ranked list of tools and subsequently select a fixed number of tools from the top (top-K) (Qu et al., 2024a; Zheng et al., 2024; Qu et al., 2024b). Although such approaches have demonstrated good performance when retrieving a single tool (Patil et al., 2023; Xu et al., 2023) or a small number of tools (generally fewer than three) (Qin et al., 2023; Huang et al., 2023), they remain susceptible to under-selection or over-selection, as illustrated in Figure.1. This limitation may prevent LLMs from addressing the current query or cause them to over-interpret the query, thereby reducing the effectiveness of LLMs in solving complex problems with external tools. Additionally, the validation of these methods often relies on datasets that use a fixed number of tools for each query, meaning that during testing, the number of tools to be used is known in advance—an unrealistic scenario in practical applications where the number of tools needed can vary dynamically. Therefore, recommending a precise and dynamically adjustable set of external tools to LLMs in a single step prior to query execution is increasingly important. This approach not only enhances the thoroughness of problem-solving but also improves efficiency by reducing the need to execute additional tools.
@@
To address these limitations, we first provide a comprehensive explanation of tool recommendation and clearly define the problem, considering the lack of definition and the incompleteness of goals pursued by existing tool retrieval methods. Toward this objective, we propose PTR, a novel model-agnostic Precision-Driven Tool Recommendation approach aimed at recommending a precise tool set for LLMs prior to query execution. By leveraging historical tool bundle usage data to uncover patterns of idiomatic use and dependencies between tools, this method is structured into three main stages: Tool Bundle Acquisition, Functional Coverage Mapping, and Multi-view-based Re-ranking. Initially, using traditional pre-trained language models, we acquire semantic matching information between queries and previously used tool bundles, thereby addressing potential performance issues of these models in zero-shot scenarios for tool recommendation tasks. Subsequently, to evaluate the effectiveness of the selected tool bundle in solving the query, LLMs are prompted to match tools with the specific subproblems they can address and to identify unresolved issues. Based on this, a multi-view-based re-ranking method is employed to select tools that can help resolve the identified issues and complement the existing tool sets. More specifically, to address the unresolved issues, we construct the final ranked list by aggregating three tool lists and ranking each tool based on their frequency of occurrence. The ranked tool list, constructed from multiple views, reduces the randomness associated with selecting tools from the entire available set.
@@
Additionally, we construct a dataset, RecTools, tailored to specific queries with recommended tool sets. In contrast to previous tool datasets that standardize the number of tools used for each query (Huang et al., 2023) or employ a small number of tools (Qu et al., 2024a), our tool recommendation set incorporates varying numbers of tools for different queries, with up to ten tools used for a single query. This is achieved through an automated process in which LLMs are prompted to generate specific queries to be addressed by given tool bundles. These queries and tool bundles are subsequently evaluated by prompting LLMs to determine whether the selected tools adequately address the corresponding queries, ensuring that neither excess nor insufficient tools are utilized. Dedicated validation and deduplication steps are implemented to ensure the precision of tool usage, thereby enhancing the quality of the tool recommendation set.
@@
Furthermore, traditional retrieval metrics such as Recall (Zhu, 2004) and Normalized Discounted Cumulative Gain (NDCG) (Järvelin & Kekäläinen, 2002), fail to capture the level of precision required for effective tool recommendation. The absence of necessary tools can lead to the failure of LLMs in performing tasks, while the redundancy of tools may cause LLMs to generate unnecessary responses. This indicates that metrics focusing solely on completeness are inadequate for evaluating tool recommendation tasks. To bridge this gap, we introduce TRACC, a novel metric designed to assess tool recommendation performance, considering both the accuracy of the quantity and the quality of the recommended tools. TRACC serves as a reliable indicator of the effectiveness of tool recommendation processes.
@@
  • We introduce tool recommendation as a novel problem, necessitating the provision of precise tool sets to LLMs for a given query. We propose PTR, an effective tool recommendation approach that leverages historical tool bundle information between queries and tools, resulting in a more accurate and comprehensive final recommended tool list.
  • We present a new dataset, RecTools, and an effective evaluation metric, TRACC, specifically designed to assess tool recommendation for LLMs. This not only addresses gaps in existing tool sets but also advances future research related to tool recommendation.
  • Extensive experiments validate the effectiveness of RecTools and demonstrate the efficacy of PTR in recommending tools for LLMs. The recommended tool sets are both comprehensive and accurate, enhancing the overall performance of LLMs in processing tasks.
2 Tool Recommendation
Tool retrieval, as discussed in previous, involves generating a comprehensive list of tools that are potentially relevant to a user’s query. This approach emphasizes breadth, aiming to maximize the inclusion of pertinent tools. While effective in ensuring extensive coverage, tool retrieval often prioritizes recall over precision, resulting in the inclusion of extraneous tools that may not be essential for the task at hand. Addressing this limitation, we propose a new optimization direction–Tool Recommendation–for LLMs. It aims to ensure that the recommended set of tools aligns closely with the ground-truth set of tools for a task, both in quantity and quality. Specifically, given a user query with a ground-truth toolset $(A,B,C)$, tool recommendation aims to identify precisely $(A,B,C)$, avoiding omissions or the inclusion of redundant tools. Here is the definition of the tool recommendation task:
Definition 1
As discussed in previous, achieving precision in tool recommendation is pivotal for enhancing the performance and reliability of LLMs. By minimizing the inclusion of irrelevant tools, LLMs can reduce computational overhead, streamline task execution, and improve the overall quality of responses. Addressing precised tool recommendation not only mitigates the drawbacks associated with broad tool retrieval but also paves the way for more sophisticated and user-centric LLM applications. This advancement is essential for deploying LLMs in environments where efficiency, accuracy, and user satisfaction are crucial.
@@ 3 The Precision-driven Tool Recommendation
We introduce a novel approach, Precision-driven Tool Recommendation (PTR), to address the challenges faced by prior research through a three-stage recommendation process: (1) Tool Bundle Acquisition, which involves establishing a potentially useful tool bundle by leveraging past usage patterns across all tool combinations, as opposed to relying solely on instructions for individual tool usage; (2) Functional Coverage Mapping, which entails effectively mapping the tools from the acquired tool bundle to the functionalities of the original query, thereby identifying which tools should be retained and which should be discarded, resulting in any remaining unsolved sub-problems; and (3) Multi-view-based Re-ranking, which involves the effective re-ranking of relevant tools from a large tool set, tailored to each unsolved sub-problem identified in the second stage, and selecting the top-ranked tool after re-ranking to complete the final recommended toolset. The overview of our approach is illustrated in Figure.2. Please note that all symbols are globally defined in sections 2 and 3. In the following sections, we present the details of these three PTR recommendation stages.
@@ 3.2 Functional Coverage Mapping
Extraction of Key Functionalities. The first step involves decomposing the user’s query $Q$ into a set of discrete and actionable functionalities $R$. This extraction ensures a comprehensive understanding of the query that the toolset must address. This extraction is achieved by prompting the language model to identify and enumerate these functionalities directly from the query, ensuring that both explicit and implicit functionalities are captured.
@@ 3.3 Multi-view based Re-ranking
Addressing the challenge of selecting pertinent tools from an extensive toolset to resolve unresolved problems requires comprehensive consideration. The proposed PTR employs a multifaceted similarity evaluation strategy that integrates three essential dimensions of the unresolved problem $U_{j}$: (1) Direct Semantic Alignment, wherein the system quantifies the semantic similarity between the user query and each available tool, ensuring the immediate identification of tools intrinsically aligned with the query’s intent; (2) Historical Query Correlation, which involves analyzing past queries that closely resemble the current one to extract tools previously utilized in similar contexts, thereby enriching the current toolset with empirically effective solutions while maintaining uniqueness through aggregation and deduplication; and (3) Contextual Tool Expansion, which leverages the most relevant tool identified through direct semantic alignment to retrieve additional tools exhibiting high similarity to this primary tool, thereby uncovering supplementary options that may offer complementary or alternative functionalities beneficial to the user’s query. The multi-view matching process involves obtaining the tool list $L$ through direct semantic alignment (DSA), historical query correlation (HQC), and contextual tool expansion (CTE), respectively. These three tool lists are then aggregated and ranked according to their frequency of occurrence, with the most frequent tools being selected. After performing the multi-view-based re-ranking for each unsolved problem, the top-ranked tool in each list is selected and added to the final recommended toolset. In some cases, it is also possible that this tool already exists in the toolset acquired from the second-stage recommendation; in such instances, the tool will be ignored. The algorithm for multi-view-based re-ranking is summarized in Algorithm.1.
@@
4:$L_{\text{DSA}}\leftarrow\text{Select}_{K}\left(\{T_{i}\in T\mid\sigma(U_{j},T_{i})\}\right)$triangleright Directly obtain the tools most relevant to the given query / /Historical Query Correlation
5:$L_{\text{HistoricalQuery}}\leftarrow\text{Select}_{K}\left(\{Q_{i}\in Q\mid\sigma(U_{j},Q_{i})\}\right)$triangleright Retrieve the most relevant past queries.
@@
13:     $T_{\text{primary}}\leftarrow L_{\text{DSA}}[0]$triangleright Obtain the most relevant tool identified in the first stage.
@@ 4 Datasets and Metrics
Table 1: Statistics of the experimental datasets.
Feature ToolLens MetaTool RecTools
Tools per Query 1-3 2 1-10
Unified used tool number
Exact Solving Test 6.34% 55.1% 61.3%
Datasets. To verify the effectiveness of PTR, we utilize three datasets for tool recommendation: ToolLens (Qu et al., 2024a), MetaTool (Huang et al., 2023), and a newly constructed dataset, RecTools. We randomly select 20% of each dataset to serve as the test data. Both ToolLens and MetaTool focus on multi-tool tasks, leading us to select them as the primary datasets for our experiments. While ToolLens uniquely emphasizes creating queries that are natural, concise, and intentionally multifaceted, MetaTool is a benchmark designed to evaluate whether LLMs possess tool usage awareness and can correctly choose appropriate tools. However, both datasets impose a low upper limit on the number of tools used per query. As the capabilities of LLMs continue to develop, more tools need to be recommended to solve increasingly complex problems, thereby limiting the applicability of these datasets. Additionally, all queries in these two datasets utilize a fixed number of tools, which not only fails to fully simulate the dynamic nature of tool usage in real-world scenarios but also introduces bias in the subsequent testing of the method. Most importantly, since tool recommendation focuses on the precision of the recommended toolset, the test datasets require that each query be exactly solvable by the provided tools (Exact Solving). Using one fewer tool leads to partial solving, while using one additional tool results in oversolving. To validate the effectiveness of the two datasets, we first employ GPT-4o as an evaluator to determine whether the provided toolset can achieve an “Exact Solving” outcome for each query. Subsequently, for each query, we randomly remove one tool from the corresponding toolset and prompt GPT-4o to assess whether the modified toolset can achieve a “Partial Solving” outcome. Queries and their respective toolsets that meet the criteria for both evaluations are considered qualified. The performance of these two datasets is not ideal. Based on these limitations, we constructed a new dataset, RecTools, where queries do not have a uniform number of tools and have a high upper limit on the number of tools used. Additionally, RecTools significantly outperforms ToolLens and Metatool in the GPT-4o “Exact Solving” test. The statistics of the three datasets are summarized in Table.1. Specifically, for all (query, tools) pairs involving the use of two and three tools, the success rates of RecTools reached 76% and 89%, respectively.
Metrics. As evaluation metrics for tool recommendation, following previous work focusing on tool retrieval (Gao et al., 2024; Qu et al., 2024b), the widely used retrieval metrics are Recall and NDCG. However, they do not adequately address the requirements for accuracy in both the number of recommended tools and the specific tools recommended, disregarding the impact of differences in size between the tool sets. Therefore, to further tailor the assessment to the challenges of tool recommendation tasks, we introduce a new metric, named TRACC. This metric is designed to measure the extent to which the recommended toolset aligns with the ground-truth set in terms of both the accuracy of the number of tools and the accuracy of the tools themselves:
@@
where $A$ denotes the ground-truth tool set and $B$ represents the recommended tool set. The cardinalities of $A$ and $B$ are denoted by $n_{1}$ and $n_{2}$, respectively. And $|A\cup B|$ signifies the cardinality of the union of $A$ and $B$. ACC represents $\frac{|A\cap B|}{n_{1}}$, where $|A\cap B|$ indicates the size of the their intersection.

### v2

# Task-Aligned Tool Recommendation for Large Language Models

@@ Abstract
By augmenting Large Language Models (LLMs) with external tools, their capacity to solve complex problems has been significantly enhanced. However, despite ongoing advancements in the parsing capabilities of LLMs, incorporating all available tools simultaneously in the prompt remains impractical due to the vast number of external tools. Consequently, it is essential to provide LLMs with a precise set of tools tailored to the specific task, considering both quantity and quality. Current tool retrieval methods primarily focus on refining the ranking list of tools and directly packaging a fixed number of top-ranked tools as the tool set. However, these approaches often fail to equip LLMs with the optimal set of tools prior to execution, since the optimal number of tools for different tasks could be different, resulting in inefficiencies such as redundant or unsuitable tools, which impede immediate access to the most relevant tools. This paper addresses the challenge of recommending precise toolsets for LLMs. We introduce the problem of tool recommendation, define its scope, and propose a novel Precision-driven Tool Recommendation (PTR) approach. PTR captures an initial, concise set of tools by leveraging historical tool bundle usage and dynamically adjusts the tool set by performing tool matching, culminating in a multi-view-based tool addition. Additionally, we present a new dataset, RecTools, and a metric, TRACC, designed to evaluate the effectiveness of tool recommendation for LLMs. We further validate our design choices through comprehensive experiments, demonstrating promising accuracy across two open benchmarks and our RecTools dataset. Our code and data are publicly available at https://github.com/GHupppp/PTR.
Task-Aligned Tool Recommendation for Large Language Models
Hang Gao Rutgers University h.gao@rutgers.edu          Yongfeng Zhang Rutgers University yongfeng.zhang@rutgers.edu
@@ 1 Introduction
Large Language Models (LLMs) have established themselves as powerful intermediaries, demonstrating remarkable impacts across a variety of downstream tasks, including text generation, code debugging, and personalized recommendations (NEURIPS2020_1457c0d6; touvron2023llama; nam2024using; chen2024large; zhao2024let). However, as these models continue to evolve, they still struggle to solve highly complex problems due to limitations arising from their pre-training data (mialon2023augmented; mallen2022not; yuan2023craft). To expand the potential of LLMs in managing more complex tasks efficiently, recommendations at various levels have been increasingly applied to LLMs. Typically, memory recommendations (borgeaud2022improving) and knowledge-based recommendations (gao2023retrieval; hu2023survey) enhance consistency and context awareness in ongoing tasks for LLMs, while data augmentation recommendations (xu2020curriculum) facilitate the inclusion of additional data to augment training. Furthermore, architecture recommendations (elsken2019neural; fedus2022switch) and prompt recommendations (shin2020autoprompt; pryzant2023automatic; liu2023pre) optimize efficiency and generate more relevant outputs. Simultaneously, to reduce the cognitive load on LLMs and enhance their complex problem-solving capabilities by enabling actions beyond natural language processing, it is crucial to augment LLMs with recommendations of optimal external tool sets, an aspect currently lacking in existing recommendation frameworks for LLMs. Furthermore, this approach will be helpful to address the challenge of input length limitations encountered when incorporating a large number of external tools into the prompt. Providing LLMs with a precise and dynamically adaptable recommended toolset can help to enhance the effectiveness of LLM’s task-solving ability.
@@
Considering that the capability of LLMs to master and control external tools is instrumental in overcoming some of their fundamental weaknesses, the field of tool retrieval, which aims to identify the top-K most suitable tools for a given query from a vast set of tools, has been increasingly explored. The advent of tool retrieval (zhuang2023toolqa; li2023api; tang2023toolalpaca; yang2024gpt4tools) signifies a nuanced evolution, most directly employing term-based methods (sparck1972statistical; robertson2009probabilistic) or semantic-based techniques (kong2023tptu; yuan2024easytool; gao2024confucius). Generally, the primary objective of these methods is to refine the ranked list of tools and subsequently select a fixed number of tools from the top (top-K) (qu2024colt; zheng2024toolrerank; qu2024tool). Although such approaches have demonstrated good performance when retrieving a single tool (patil2023gorilla; xu2023tool) or a small number of tools (generally fewer than three) (qin2023toolllm; huang2023metatool), they remain susceptible to under-selection or over-selection, as illustrated in Figure.1. This limitation may prevent LLMs from addressing the current query or cause them to over-interpret the query, thereby reducing the effectiveness of LLMs in solving complex problems with external tools. Additionally, the validation of these methods often relies on datasets that use a fixed number of tools for each query, meaning that during testing, the number of tools to be used is known in advance—an unrealistic scenario in practical applications where the number of tools needed can vary dynamically. Therefore, recommending a precise and dynamically adjustable set of external tools to LLMs in a single step prior to query execution is increasingly important. This approach not only enhances the thoroughness of problem-solving but also improves efficiency by reducing the need to execute additional tools.
@@
To address these limitations, we first provide a comprehensive explanation of tool recommendation and clearly define the problem (Appendix A), considering the lack of definition and the incompleteness of goals pursued by existing tool retrieval methods. Toward this objective, we propose PTR, a novel model-agnostic Precision-Driven Tool Recommendation approach aimed at recommending a precise tool set for LLMs prior to query execution. By leveraging historical tool bundle usage data to uncover patterns of idiomatic use and dependencies between tools, this method is structured into three main stages: Tool Bundle Acquisition, Functional Coverage Mapping, and Multi-view-based Re-ranking. Initially, using traditional pre-trained language models, we acquire semantic matching information between queries and previously used tool bundles, thereby addressing potential performance issues of these models in zero-shot scenarios for tool recommendation tasks. Subsequently, to evaluate the effectiveness of the selected tool bundle in solving the query, LLMs are prompted to match tools with the specific subproblems they can address and to identify unresolved issues. Based on this, a multi-view-based re-ranking method is employed to select tools that can help resolve the identified issues and complement the existing tool sets. More specifically, to address the unresolved issues, we construct the final ranked list by aggregating three tool lists and ranking each tool based on their frequency of occurrence. The ranked tool list, constructed from multiple views, reduces the randomness associated with selecting tools from the entire available set.
@@
Additionally, we construct a dataset, RecTools, tailored to specific queries with recommended tool sets. In contrast to previous tool datasets that standardize the number of tools used for each query (huang2023metatool) or employ a small number of tools (qu2024colt), our tool recommendation set incorporates varying numbers of tools for different queries, with up to ten tools used for a single query. This is achieved through an automated process in which LLMs are prompted to generate specific queries to be addressed by given tool bundles. These queries and tool bundles are subsequently evaluated by prompting LLMs to determine whether the selected tools adequately address the corresponding queries, ensuring that neither excess nor insufficient tools are utilized. Dedicated validation and deduplication steps are implemented to ensure the precision of tool usage, thereby enhancing the quality of the tool recommendation set.
@@
Furthermore, traditional retrieval metrics such as Recall (zhu2004recall) and Normalized Discounted Cumulative Gain (NDCG) (jarvelin2002cumulated), fail to capture the level of precision required for effective tool recommendation. The absence of necessary tools can lead to the failure of LLMs in performing tasks, while the redundancy of tools may cause LLMs to generate unnecessary responses. This indicates that metrics focusing solely on completeness are inadequate for evaluating tool recommendation tasks. To bridge this gap, we introduce TRACC, a novel metric designed to assess tool recommendation performance, considering both the accuracy of the quantity and the quality of the recommended tools. TRACC serves as a reliable indicator of the effectiveness of tool recommendation processes.
@@
1. We introduce tool recommendation as a novel problem, necessitating the provision of precise tool sets to LLMs for a given query. We propose PTR, an effective tool recommendation approach that leverages historical tool bundle information between queries and tools, resulting in a more accurate and comprehensive final recommended tool list.
2. We present a new dataset, RecTools, and an effective evaluation metric, TRACC, specifically designed to assess tool recommendation for LLMs. This not only addresses gaps in existing tool sets but also advances future research related to tool recommendation.
3. Extensive experiments validate the effectiveness of RecTools and demonstrate the efficacy of PTR in recommending tools for LLMs. The recommended tool sets are both comprehensive and accurate, enhancing the overall performance of LLMs in processing tasks.
@@ 2 The Precision-driven Tool Recommendation
We introduce a novel approach, Precision-driven Tool Recommendation (PTR), to address the challenges faced by prior research through a three-stage recommendation process: (1) Tool Bundle Acquisition, which involves establishing a potentially useful tool bundle by leveraging past usage patterns across all tool combinations, as opposed to relying solely on instructions for individual tool usage; (2) Functional Coverage Mapping, which entails effectively mapping the tools from the acquired tool bundle to the functionalities of the original query, thereby identifying which tools should be retained and which should be discarded, resulting in any remaining unsolved sub-problems; and (3) Multi-view Based Re-ranking, which involves the effective re-ranking of relevant tools from a large tool set, tailored to each unsolved sub-problem identified in the second stage, and selecting the top-ranked tool after re-ranking to complete the final recommended toolset. The overview of our approach is illustrated in Figure.2. Please note that all symbols are globally defined in Section 2. In the following sections, we present the details of these three PTR recommendation stages.
@@ 2.2 Functional Coverage Mapping
Extraction of Key Functionalities. The first step involves decomposing the user’s query $Q$ into a set of discrete and actionable functionalities $F$. This extraction ensures a comprehensive understanding of the query that the toolset must address. This extraction is achieved by prompting the language model to identify and enumerate these functionalities directly from the query, ensuring that both explicit and implicit functionalities are captured.
@@ 2.3 Multi-view Based Re-ranking
Addressing the challenge of selecting pertinent tools from an extensive toolset to resolve unresolved problems requires comprehensive consideration. The proposed PTR employs a multifaceted similarity evaluation strategy that integrates three essential dimensions of the unresolved problem $U_{j}$: (1) Direct Semantic Alignment, wherein the system quantifies the semantic similarity between the user query and each available tool, ensuring the immediate identification of tools intrinsically aligned with the query’s intent; (2) Historical Query Correlation, which involves analyzing past queries that closely resemble the current one to extract tools previously utilized in similar contexts, thereby enriching the current toolset with empirically effective solutions while maintaining uniqueness through aggregation and deduplication; and (3) Contextual Tool Expansion, which leverages the most relevant tool identified through direct semantic alignment to retrieve additional tools exhibiting high similarity to this primary tool, thereby uncovering supplementary options that may offer complementary or alternative functionalities beneficial to the user’s query. The multi-view matching process involves obtaining the tool list $L$ through direct semantic alignment (DSA), historical query correlation (HQC), and contextual tool expansion (CTE), respectively. These three tool lists are then aggregated and ranked according to their frequency of occurrence, with the most frequent tools being selected. After performing the multi-view based re-ranking for each unsolved problem, the top-ranked tool in each list is selected and added to the final recommended toolset. In some cases, it is also possible that this tool already exists in the toolset acquired from the second-stage recommendation; in such instances, the tool will be ignored. The algorithm for multi-view-based re-ranking is summarized in Algorithm.1.
@@
[…]
@@
13:  $T_{\text{primary}}\leftarrow L_{\text{DSA}}[0]$
@@ 3 Datasets and Metrics
Datasets. To verify the effectiveness of PTR, we utilize three datasets for tool recommendation: ToolLens (qu2024colt), MetaTool (huang2023metatool), and a newly constructed dataset, RecTools. We randomly select 20% of each dataset to serve as the test data. Both ToolLens and MetaTool focus on multi-tool tasks, leading us to select them as the primary datasets for our experiments. While ToolLens uniquely emphasizes creating queries that are natural, concise, and intentionally multifaceted, MetaTool is a benchmark designed to evaluate whether LLMs possess tool usage awareness and can correctly choose appropriate tools.
Metrics. As evaluation metrics for tool recommendation, following previous work focusing on tool retrieval (gao2024confucius; qu2024tool), the widely used retrieval metrics are Recall and NDCG. However, they do not adequately address the requirements for accuracy in both the number of recommended tools and the specific tools recommended, disregarding the impact of differences in size between the tool sets. Therefore, to further tailor the assessment to the challenges of tool recommendation tasks, we introduce a new metric, named TRACC. This metric is designed to measure the extent to which the recommended toolset aligns with the ground-truth set in terms of both the accuracy of the number of tools and the accuracy of the tools themselves:
@@
where $A$ denotes the ground-truth tool set and $B$ represents the recommended tool set. The cardinalities of $A$ and $B$ are denoted by $n_{1}$ and $n_{2}$, respectively. And $|A\cup B|$ signifies the cardinality of the union of $A$ and $B$. ACC represents $\frac{|A\cap B|}{n_{1}}$, where $|A\cap B|$ indicates the size of the their intersection. More details can be found in Appendix C.